OpenStudy (anonymous):

Jamie has a deck of 60 sports cards, of which some are baseball cards and some are football cards. Jamie pulls out a card randomly from the deck, records its type, and replaces it in the deck. Jamie has already recorded three baseball cards and nine football cards. Based on these data, what is, most likely, the number of baseball cards in the deck? 12 15 24 30

OpenStudy (anonymous):

i believe it is 24

OpenStudy (tkhunny):

Proportion: \(Drawn\;vs.\;Deck \longrightarrow \dfrac{3\; Baseball}{9\;Football} = \dfrac{Total\; Baseball}{Total\;Football}\) Further, Total Baseball + Total Football = 60 How did you get 24? I'm not seeing that as the answer.

OpenStudy (anonymous):

oh so it is 15

OpenStudy (tkhunny):

Please learn to show your work and communicate your thinking. It is 15.

OpenStudy (anonymous):

3/12 is = to 1/4 so 1/4 of 60=15

OpenStudy (anonymous):

that is how i solved it